### v1

# NFIG: Autoregressive Image Generation with Next-Frequency Prediction

Abstract
Autoregressive models have achieved promising results in natural language processing. However, for image generation tasks, they encounter substantial challenges in effectively capturing long-range dependencies, managing computational costs, and most crucially, defining meaningful autoregressive sequences that reflect natural image hierarchies. To address these issues, we present Next-Frequency Image Generation (NFIG), a novel framework that decomposes the image generation process into multiple frequency-guided stages. Our approach first generates low-frequency components to establish global structure with fewer tokens, then progressively adds higher-frequency details, following the natural spectral hierarchy of images. This principled autoregressive sequence not only improves the quality of generated images by better capturing true causal relationships between image components, but also significantly reduces computational overhead during inference. Extensive experiments demonstrate that NFIG achieves state-of-the-art performance with fewer steps, offering a more efficient solution for image generation, with 1.25 speedup compared to VAR-d20 while achieving better performance (FID: 2.81) on the ImageNet-256 benchmark. We hope that our insight of incorporating frequency-domain knowledge to guide autoregressive sequence design will shed light on future research. We will make our code publicly available upon acceptance of the paper.
1 Introduction
The synthesis of realistic and diverse images has emerged as a fundamental challenge in computer vision. Rapid progress in this field has been propelled by the success of deep generative models, such as generative adversarial networks [6, 12, 32, 13], diffusion models [8, 41, 7, 39], and autoregressive models (AR) [35, 10, 21, 38].
Despite remarkable advances in existing methods, AR models for image generation still face several fundamental challenges.
On the one hand, due to their inherently local and sequential nature, most current AR models struggle to effectively capture long-range dependencies and global structure [3]. For example, PixelCNN [35] generates an image by predicting each pixel in a raster scanning sequence, which neglects the global image structure and relationships with distant elements.
On the other hand, the generation process is computationally intensive and time-consuming, as AR models always generate pixels or patches sequentially in a predetermined order, with each new element requiring the computation of conditional probabilities based on all previously generated content [22, 33]. For instance, ViTVQ [40] requires more than 6 seconds to generate a image over 1024 steps, making it impractical for real-time applications.
Most importantly, AR models face a fundamental challenge in defining a meaningful autoregressive sequence. Traditional AR models using raster scanning or predefined arbitrary orders fail to reflect the natural hierarchical structure and dependencies in images[34]. This improper sequence design makes it difficult for models to capture the true causal relationships between different image components, ultimately affecting the coherence and visual quality of generated outputs.
To address these limitations, recent works have explored incorporating various improvements into the generation process.
For example, Taming Transformer [5] partially addresses long-range dependency challenges through its discrete latent space and transformer architecture, but still suffers from computational inefficiency. Fast PixelCNN++ [27] speeds up generation in convolutional autoregressive models by caching hidden states to avoid redundant computation, achieving up to 183× speedups, yet it doesn’t fundamentally change the autoregressive sequence design. VAR [34] leverages the Laplacian Pyramid as a prior to guide autoregressive image generation across different resolutions, achieving improved generation quality with reduced computational overhead. However, these methods do not fully exploit the potential of natural priors inherent in raw images to guide the generation process and improve the efficiency of AR models.
In fact, the natural structure of images follows a hierarchical frequency distribution—low frequencies encode global structures while high frequencies contain local details. This organization suggests an efficient autoregressive generation sequence from low to high frequencies, aligning with visual information’s natural structure. Since low-frequency components require fewer tokens to represent, this approach enhances computational efficiency. Similar frequency-progressive principles have proven effective in diffusion models, which build from low-frequency foundations before adding higher-frequency details [26].
Building on this insight, we propose a Next-Frequency Image Generation (NFIG) framework for AR models that: (1) first generates a low-frequency image with few tokens to capture global structure; (2) then progressively adds higher-frequency components conditioned on the low-frequency foundation. This process has been shown in Figure 1. Grounded in information theory, this approach efficiently represents information across the frequency spectrum using our Frequency-guided Residual-quantized VAE (FR-VAE).
Our main contributions can be summarized as three points:
We introduce a Next-Frequency Image Generation (NFIG) framework that incorporates frequency analysis into AR image generation. To the best of our knowledge, this is the first work that guides the autoregressive process by leveraging the frequency spectrum of the image;
To demonstrate the feasibility of the NFIG paradigm, we design a Frequency-guided Residual-quantized VAE as our image tokenizer. FR-VAE separates low and high-frequency components in the representation learning process, with low frequencies encoding global structure and high frequencies preserving local details. Experiments show FR-VAE achieves a reconstruction FID of 0.85, validating its image content preservation capability;
Extensive experiments show that our proposed method achieves state-of-the-art image generation quality, reaching an FID of 2.81 with a relatively small model size. This improvement paves the way for more effective and efficient AR image generation models, making them more practical for real-world applications.
2 Related Work
2.1 Image Tokenizer
Image tokenizers, which transform continuous image data into discrete representations, have become a critical component in modern image generation systems. VQ-VAE [36] introduces vector quantization into VAE, reducing the pressure on the downstream generative model by transforming the continuous latent space into a discrete one. VQ-VAE-2 [28] extends this idea with a hierarchical framework and multi-scale codebooks, where top-level codes capture global structure, and bottom-level codes model local details, enabling higher-quality reconstruction and generation at increased resolutions.
However, VQ-VAE-based methods often suffer from codebook collapse, where only a few codebook entries are effectively used. FSQ [23] attempts to address this issue by utilizing finite scalar quantization to learn the codebook, but it does not learn a meaningful feature of images.
RQ-VAE [14] employs residual quantization with a shared codebook to enhance reconstructed image quality.
XQGAN [19] introduces feature product decomposition and a residual quantizer to enhance VQ-VAE’s performance, leading to improved image generation results.
While these approaches have made significant progress in image tokenization, they often neglect the inherent multi-scale structure of natural images, which is crucial for efficient and effective representation learning.
2.2 Autoregressive Image Generation
Autoregressive image generation has demonstrated remarkable capabilities in producing high-quality images by modeling the joint distribution of image tokens as a product of conditionals[20]. PixelCNN [35] generates images pixel by pixel in a sequential order (typically from top-left to bottom-right) using masked convolutions to ensure each pixel is conditioned only on previously generated pixels. Taming Transformer [5] introduces an autoregressive approach that generates high-resolution images by predicting the next latent patch token in a discrete compressed space learned through vector quantization. Emu3 [37] patchifies an image into a series of tokens and generates images by predicting tokens in a raster-scan manner. VAR [34] incorporates the prior knowledge of Laplacian Pyramid into transformer architecture and generates images in a next-resolution manner. MAR [17] improves the quality of generated images by replacing discrete tokens with continuous features, recognizing that autoregressive models primarily need per-token probability distributions. Infinity [17] combines autoregressive modeling with Bit-wise Modeling to enhance visual details in high-resolution image synthesis. ImageFolder [18] utilizes folded image tokens to generate high-quality images in a next-scale prediction manner, achieving superior performance.
These approaches collectively demonstrate the evolution of autoregressive image generation techniques, progressing from pixel-level prediction to more sophisticated methods involving latent spaces, hierarchical structures, and physical priors. Despite their differences in implementation, all these methods share the fundamental autoregressive principle of sequentially generating image elements conditioned on previously generated content.
3 Method
Figure 2 illustrates the main workflow for NFIG, which generates an image from low-frequency to high-frequency components, iteratively improving image quality.
3.1 Frequency-guided Residual-quantized VAE
To generate images in a frequency-aware manner, we propose a Frequency-guided Residual-quantized VAE (FR-VAE) with VQ-GAN framework. The key idea is to represent lower-frequency signals with fewer tokens and higher-frequency components with more tokens. The workflow of FR-VAE has been shown in Figure 3.
3.1.1 Frequency-guided Reconstruction
As shown in Figure 4, raw images can be decomposed into components across frequency bands: low frequencies capture global structure, high frequencies preserve details. Using Frequency-guided Decomposer and Composer, these components can be losslessly recombined for complete visual representation.
Frequency-guided Decomposer.
Given a image and a encoder , there is image latent feature and . FR-VAE decomposes into several component with different frequency by Frequency-guided Decomposer via Fast Fourier Transform (FFT) as follows:
where is the element-wise product, denotes the FFT operation, is the inverse FFT, and is the -th frequency mask used to select the desired frequency range, is the total number of frequency masks, is the component corresponding to .
Frequency-guided Composer.
Frequency-guided Composer reconstructs the raw image by interpolating different frequency components to a uniform size and merging them into a single image, as illustrated in Figure 4.
where is the interpolation function, which enables the Frequency-guided Composer to process images of different frequencies at varying resolutions.
3.1.2 Frequency-guided Residual-quantization
To efficiently represent images with minimal tokens, we implement a frequency-guided residual quantization approach that addresses information loss during downsampling. Our method progressively captures different frequency components of an image through a residual learning scheme.
Residual Token Extraction.
Given a sequence of feature maps with different dimensions , where and if , and and , we supervise the learning process using accumulated signals from the lowest frequency to the current frequency band.
The residual and representation of the -th frequency component can be computed as follows:
where is the interpolation function that upsamples
to the original feature map size, and represents the difference between the accumulated frequency components up to the -th level and the learnable features.
Vector Quantization.
In general, autoregressive models utilize the discrete tokens to generate a image. To achieve this goal, we take a simple vector quantization to transform the continuous token into discrete tokens.
We define a quantizer with a learnable codebook containing code vectors.
Using this codebook, the quantizer transforms a continuous feature map into a set of discrete tokens , where each token has the corresponds to a vector . The process of finding the optimal code representation involves:
where lookup is a function that take the index of from codebook . Due to the difficulty in learning an effective codebook, we have introduced the DINO model as an adversarial discriminator to improve codebook .
3.2 Autoregressive Image Generation
To generate images progressively from low to high frequency components, we implement a decoder-only transformer framework with Adaptive Layer Normalization (AdaLN) and block-wise causal attention [34].
3.2.1 Next-Frequency Image Prediction
Unlike conventional autoregressive image generation models that employ a ”token-by-token prediction” strategy, which often neglects spatial relationships and inherent image structure. NFIG adopts a “Coarse-to-Fine Generation” approach, first synthesizes the low-frequency components of an image, then iteratively incorporates higher-frequency details, progressively refining the generated output at each step.
The generation process for next-frequency prediction can be formulated as:
where represents the matrix of code indices utilized to encode the -th frequency component, and denotes the set of possible indices.
3.2.2 Frequency Band Division Strategy
We treat the lower frequency components as the foundation or “basis” for generating , represented by . According to information theory principles, lower-frequency signals contain less information and require fewer tokens, while higher-frequency components carry more detailed information and need more tokens for accurate representation.
Consequently, we establish an increasing scale sequence for components with increasing frequency bands . Here, denotes the maximum frequency of the entire image feature map , with . We divide the frequency bands based on their corresponding resolution as:
This frequency-guided progressive approach allows our model to capture and prioritize salient components at each stage. The method improves both computational efficiency and image quality by explicitly modeling the multi-scale frequency structure inherent in natural images.
| Type | Model | rFID | gFID | IS | Pre | Rec | #Para | #Step | Time |
| GAN | BigGAN [2] | - | 6.95 | 224.5 | 0.89 | 0.38 | 112M | 1 | – |
| GAN | GigaGAN [11] | - | 3.45 | 225.5 | 0.84 | 0.61 | 569M | 1 | – |
| GAN | StyleGan-XL [32] | - | 2.30 | 265.1 | 0.78 | 0.53 | 166M | 1 | 0.75 |
| Diff. | ADM [4] | - | 10.94 | 101.0 | 0.69 | 0.63 | 554M | 250 | 420 |
| Diff. | CDM [9] | - | 4.88 | 158.7 | – | – | – | 8100 | – |
| Diff. | LDM-4-G [30] | - | 3.60 | 247.7 | – | – | 400M | 250 | – |
| Diff. | DiT-L/2 [25] | 0.9 | 5.02 | 167.2 | 0.75 | 0.57 | 458M | 250 | 77.5 |
| Diff. | DiT-XL/2 [25] | 0.9 | 2.27 | 278.2 | 0.83 | 0.57 | 675M | 250 | 112.5 |
| Diff. | L-DiT-3B [1] | 0.9 | 2.10 | 304.4 | 0.82 | 0.60 | 3.0B | 250 | 112.5 |
| Diff. | L-DiT-7B [1] | 0.9 | 2.28 | 316.2 | 0.83 | 0.58 | 7.0B | 250 | 112.5 |
| Mask. | MaskGIT [29] | 2.28 | 6.18 | 182.1 | 0.80 | 0.51 | 227M | 8 | 1.25 |
| Mask. | RCG [16] | - | 3.49 | 215.5 | – | – | 502M | 20 | 4.75 |
| AR | VQVAE-2 [29] | 2.0 | 31.11 | 45.0 | 0.36 | 0.57 | 13.5B | 5120 | – |
| AR | VQGAN [5] | 7.94 | 18.65 | 80.4 | 0.78 | 0.26 | 227M | 256 | 47.5 |
| AR | VQGAN [5] | 7.94 | 15.78 | 74.3 | – | – | 1.4B | 256 | 60 |
| AR | ViTVQ [40] | 1.28 | 4.17 | 175.1 | – | – | 1.7B | 1024 | 60 |
| AR | ViTVQ-re [40] | 1.28 | 3.04 | 227.4 | – | – | 1.7B | 1024 | 60 |
| AR | RQTran. [15] | 1.83 | 7.55 | 134.0 | – | – | 3.8B | 68 | 52.5 |
| AR | RQTran.-re [15] | 1.83 | 3.80 | 323.7 | – | – | 3.8B | 68 | 52.5 |
| AR | VAR-d16 [34] | 0.9 | 3.55 | 274.4 | 0.84 | 0.51 | 310M | 10 | 1 |
| AR | VAR-d20 [34] | 0.9 | 2.95 | 302.6 | 0.83 | 0.56 | 600M | 10 | 1.25 |
| AR | NFIG(Ours) | 0.85 | 2.81 | 332.42 | 0.77 | 0.59 | 310M | 10 | 1 |
4 Experiment
This section introduces the experimental setup, including datasets, metrics, baselines, and implementation details. We then compare NFIG with state-of-the-art approaches on image generation tasks, followed by ablation studies and motivation verification experiments to analyze component effectiveness and design choices.
4.1 Experimental Settings
Dataset. For the purpose of our experiments, we use the ILSVRC 2012 subset of ImageNet [31], which contains 1.2 million training images, 50,000 validation images, and 100,000 test images. This subset focuses on 1,000 object categories, with each category having approximately 1,200 training images, 50 validation images, and 100 test images.
Evaluation Metrics. We adopt four metrics for quantitative evaluation: Fréchet Inception Distance (FID), Inception Score (IS), Precision (Pre), and Recall (Rec).
Baselines. We compare our method with several state-of-the-art image generation approaches, including generative adversarial networks (GAN.), diffusion models (Diff.), mask diffusion (Mask.), and autoregressive model (AR.). These methods have achieved impressive results on various image generation tasks and serve as strong baselines for comparison.
Implementation Details. We implement our model using PyTorch [24] and train it on NVIDIA H100 GPUs. For the image tokenizer, FR-VAE follows the VQGAN architecture with DINO as discriminator. The image encoder leverages pretrained weights from DINOv2-base. Our frequency residual quantizer employs multiple scaling factors across different frequency bands, resulting in a total of 680 tokens. We use a codebook size of 4096 for FR-VAE. The image generator’s backbone is based on the VAR Transformer architecture with depth 16, which enables multi-scale image prediction. We optimize our model using the Adam optimizer with a learning rate of 8e-5 and a batch size of 768. The model is trained for 350 epochs on the ImageNet dataset. At the inference stage, we set CFG 4.5 and top_k 990.
4.2 Main Results
Table 1 presents a comprehensive comparison between our approach and state-of-the-art image generative models on ImageNet . The results demonstrate that NFIG achieves superior performance within the AR model family while establishing itself as a formidable competitor among diverse generative methods across different paradigms.
AR Model Comparison. NFIG achieves the best gFID (2.81) and IS (332.42) scores, significantly outperforming other AR models. Compared to VAR-16 (gFID: 3.55, IS: 274.4), our approach reduces FID by 0.74 and improves IS by more than 21%. Additionally, the proposed approach outperforms VAR-d20, a relatively larger model, while delivering 25% faster inference speed.
Cross-family Comparison. NFIG achieves competitive performance with the best models from other families. NFIG outperforms the best mask diffusion model RCG, which has a gFID score of 3.49 and IS score of 215.5. While some GANs like StyleGAN-XL have lower gFID scores (2.30) with moderate IS (265.1), and diffusion models like DiT-L/2 show excellent gFID (2.27) and strong IS (278.2), NFIG uniquely balances both metrics at high levels (gFID: 2.81, IS: 332.42). This establishes NFIG as not only the leading AR model but also a strong competitor across all model types.
Qualitative Results. The qualitative visualization of NFIG on ImageNet is shown in Figure 5. Our model demonstrates remarkable generation capabilities across diverse ImageNet categories.
4.3 Ablation Study
To investigate the effectiveness of different components in our proposed NFIG model, we conduct a thorough ablation study on the ImageNet validation set. The results are presented in Table 2, where we evaluate the performance using rFID and gFID metrics.
We start with the baseline AR model with a sequence length of 256, which achieves an rFID of 1.62 and a gFID of 18.65.
Image Tokenizer.
To further enhance the model’s performance, we progressively introduce additional components. First, incorporating frequency-guided residual quantization into the VAR framework reduces the rFID to 1.40. Next, we integrate the DINO discriminator from VAR’s tokenizer, which substantially improves the rFID from 1.40 to 0.85.
Transformer.
We then utilize the image tokenizer (FR-VAE) to train the transformer model. Without Top_k and Classifier Free Guidance (CFG), NFIG achieves a gFID of 9.7.
The addition of Top_k sampling strategy further reduces the gFID to 6.83. Finally, incorporating CFG yields the best overall performance, maintaining the rFID at 0.85 while dramatically improving the gFID to 2.81.
Our experimental results demonstrate that the combination of FR-VAE with CFG provides optimal generation quality.
Moreover, we observe that both DINO discriminator and FR-VAE contributes significantly to improving rFID for the image tokenizer. Additionally, Top_k sampling and CFG prove essential for reducing gFID. These findings highlight the importance of discriminator guidance and conditional generation strategies in enhancing image generation quality.
| type | Method | Length | Evaluation Metric | |
| rFID | gFID | |||
| 1 | AR | 256 | 1.62 | 18.65 |
| Image Tokenizer | ||||
| 2 | + FR-Quantizer | 680 | 1.40 | – |
| 3 | + DINO-Disc | 680 | 0.85 | – |
| Generation Transformer | ||||
| 4 | + AdaLN | 680 | 0.85 | 9.7 |
| 5 | + Top_k | 680 | 0.85 | 6.83 |
| 6 | + CFG | 680 | 0.85 | 2.81 |
4.4 Motivation Verification
Effectiveness of Frequency Guidance. Similar to NFIG, VAR follows a ”coarse-to-fine” approach but differs significantly in loss computation across resolutions. VAR computes loss between different resolutions and the raw image, causing disproportionately large loss values at lower resolutions. In contrast, NFIG utilizes frequency components to guide feature learning, providing more balanced loss values throughout generation. As Figure 7 shows, this leads to dramatic variations in vector quantization loss—VAR exhibits substantially higher values across all scale factors, while NFIG maintains considerably lower loss values across all resolutions.
Frequency Distribution Analysis. The experimental results in Figure 6 demonstrate the progressive refinement of generated images and the effective capture and synthesis of multi-scale visual features by the NFIG model. The frequency spectrum visualizations reveal the model’s ability to hierarchically incorporate information from low to high frequencies, resulting in generated images with rich details and natural appearance.
5 Conclusion
This paper introduces Next-Frequency Image Generation, a novel autoregressive framework that decomposes image generation into frequency-guided stages. Our key insight leverages the hierarchical spectral distribution of natural images: low-frequency components encode global structures and long-range dependencies, while high-frequency components contain local details requiring greater information entropy. By progressively generating from low to high frequencies, the proposed method significantly outperforms existing models with comparable parameter counts, demonstrating superior quality metrics while maintaining computational efficiency. Experiments confirm that our frequency-guided approach represents an important advancement in autoregressive image synthesis.